OpenStudy (hitaro9):

Let a be a positive integer. Find all positive remainders when a^100 is divided by 125

OpenStudy (hitaro9):

So just by plugging in numbers in to wolfram alpha, it looks like the remainder is going to be 0 (for 5^n as a) and 1 (for all other numbers). I'm having trouble seeing why though.

ganeshie8 (ganeshie8):

Familiar with Euler generalization to Fermat's little theorem ?

OpenStudy (hitaro9):

Yes.

ganeshie8 (ganeshie8):

then we're done, simply use it :)

ganeshie8 (ganeshie8):

when \(a=5k\) we have \[a^{100}= (5k)^{100} \equiv 0 \pmod{125}\] when \(\gcd(a, 125) = 1\), by euler generelization to fermat's little thm we have \[a^{\phi(125)}\equiv 1 \pmod{125}\]

ganeshie8 (ganeshie8):

not surprisingly \(\phi(125) = \phi(5^3) = 5^3 - 5^2 = 100\)

OpenStudy (hitaro9):

Oh right. That makes sense. Lot simpler than I was thinking thank you.

ganeshie8 (ganeshie8):

np:)